OpenStudy (anonymous):

What is the solution of log3x − 2 125 = 3?

OpenStudy (anonymous):

the 3x-2 is small.

OpenStudy (anonymous):

\[\log(3x-2) +125=3?\]

OpenStudy (anonymous):

yup. but the 3x-2 is small

OpenStudy (anonymous):

\[\log_{3x-2}125=3 \] like this?

OpenStudy (anonymous):

This is quite simple. That is the same as writing \[(3x-2)^{3}=125\]

OpenStudy (anonymous):

we can now expand and simplify and solve for x

OpenStudy (anonymous):

actually a more efficient way of doing this is by taking the cube root of both sides\[3x-2=5\] \[x=\frac{ 7 }{ 3 }\]

OpenStudy (anonymous):

Thanks! I have another question though: The frog population at a lake doubles every week. The population can be modeled by f(x) = 15(2)x and f(6) = 960. What does the 6 represent? The population after six weeks The starting population The rate the population increases The number of weeks that have passed